how can we compare coloumbs inverse square law in gravitation with electrostatics?
 

Grade:12th pass

2 Answers

Vallakatla Aditya
4 years ago
The coloumbs inverse square law in gravitation and electrostatics is similar.
In gravitation the formula is Gm1m2/r2 , now in electrostatics the formula is Kq1q2/r here K=1/4piE ,
E= permittivity in free space .       In those exprssions G is replaced by K and m is replaced by q .
Khimraj
4 years ago
  1. People forget that the electric force is inversely proportional to r^2r2r, start superscript, 2, end superscriptinstead of rrr. As the distance between charges increases, the electric force decreases by a factor of \dfrac{1}{r^2}r21start fraction, 1, divided by, r, start superscript, 2, end superscript, end fraction. For example, if we double the distance between the two electrons, the repulsive force between them would reduce (because it is inverse), and it would go down by a factor of 444instead of 222 (because of the square).
  2. People get confused about finding the sign of F_EFEF, start subscript, E, end subscript. Notice that our equation for F_EFEF, start subscript, E, end subscript uses absolute values for the force and charge. Instead of calculating the force direction of F_EFEF, start subscript, E, end subscript by inputting the charge signs for its equation, we determine the direction by considering what combination of charges we have. As seen in Figures 1 and 2, opposite charges attract each other with force F_EFEF, start subscript, E, end subscript and like charges repel each other with F_EFEF, start subscript, E, end subscript.